# Multi–charged geometries with cosmological constant

### v1

@@ 1 Introduction
Over the last three decades, studies of black holes and black branes have led to remarkable progress in understanding string theory [1, 2], quantum gravity [3] and strongly coupled systems [4, 5, 6]. In particular, an important role has been played by charged objects that generalize the classic cases of the Reissner–Nordström and Kerr–Newman black holes in several directions: they contain extended sources [1], multiple gauge fields [7], nontrivial scalars [8, 9], and the cosmological constant [10]. While many systems without the cosmological constant arise from the ten–dimensional supergravities and their reductions, one can also extend these constructions to an arbitrary number of dimensions by starting from the Myers–Perry black holes [11] and dressing them with charges using techniques developed in [7]. In the presence of the cosmological constant, the neutral solutions have also been constructed in arbitrary dimensions [10], but addition of charge presents significant challenges, and has only been done on the case–by–case basis. Such charged extensions were mostly constructed in the context of gauged supergravities, which are defined only in some dimensions, and whose field contents vary significantly as one looks at different dimensions. In this article we will use known solutions of several gauged supergravities as inspirations for constructing asymptotically-AdS black holes with an arbitrary number of charges in an arbitrary number of dimensions. We will also construct closely related geometries produced by $p$–branes, extending the well-known results obtained in [1].
@@
The great success in constructing charged black holes and black branes in string theory is partly attributed to the discovery of the U–duality group that allows one to start with a neutral seed solution of nonlinear differential equations and dress it with charges using a pure algebraic procedure [^3] The discovery of black branes in supergravity by Horowitz and Strominger [1] pre-dates this algebraic method. Furthermore, the algebraic procedure applies only for specific couplings between gauge fields and scalars, and some solutions found in [1] and related work [8] come from more general Lagrangians.. At the level of classical gravity for bosonic fields, such U–duality group can be formally defined in any number of dimensions, allowing one to start from a product of the neutral Myers–Perry geometry [11] and a torus and add several charges to the solutions. In the process of doing so, one generates gauge fields as well as several scalars. Since the reduction is performed on a torus, the gauge fields are abelian, and scalars don’t have a potential. A similar reduction on spheres is more complicated, and often it leads to gauged supergravities with a nontrivial potential for the scalars. Unfortunately, such reduction does not produce a U–duality group large enough to algebraically generate new charged solutions from the neutral ones, so the relevant configurations have been found on the case–by–case basis by solving the equations of motion. Furthermore, the actions for gauged supergravities and even their field contents strongly depend on dimensions. On the other hand, reduction on spheres also has a significant advantage: due to a nontrivial potential, the reduced theory admits solutions with anti–de–Sitter asymptotics, while the geometries obtained from the toric case are generally asymptotically–flat [^4] One can also construct solutions with other asymptotics by taking various limits, but if all charges are turned off, then the geometries obtained via toric reduction always approach flat space at infinity.. Therefore, the geometries constructed in gauged supergravities have played a prominent role in the exploration of the AdS/CFT correspondence [5, 6], and it is very interesting to find new examples of such solutions. In this article we will use solutions of gauged supergravity as inspirations for a general ansatz describing asymptotically AdS black holes with an arbitrary number of charges and find the unique action, which depends only on the numbers of fields and dimensions, that gives rise to such solutions. All known gauged supergravities are reproduced as special cases.
@@
While many black hole geometries with AdS asymptotics are known, finding black branes with such asymptotics presents a significant challenge, even in the neutral case. This problem can be attributed to a relatively small amount of symmetry in the resulting configuration: while black holes are naturally embedded in the AdS space in the standard coordinates, black branes violate many symmetries of this space. Therefore, the ansatz interpolating between black branes and AdS asymptotics contains many degrees of freedom, and all attempts to obtain the solutions of equations of motion in closed analytical form have been unsuccessful [^5] This discussion focuses on static flat branes. The geometries produced by supersymmetric branes with curved worldvolumes have been constructed [12], but they contain rotation, as expected from the general properties of such branes [13].. The study of extended black objects in the presence of the cosmological constant began with exploration of black strings with AdS5 asymptotics [14], and it was later extended to black strings in arbitrary dimensions in [15, 16, 17]. The authors of these articles extracted many important properties of black strings with cosmological constant, but unfortunately, they had to rely on numerical simulations. More recent article [18] extended these results to arbitrary $p$–branes and complemented the numerical analysis of [14, 15] by various analytical expansions, and in some special cases it reduced the problem to solutions of a relatively simple Abel equation. All these analyses were performed for neutral strings/branes, and in this article we extend them to the charged case. We will use the analytical expansions introduced in [18] to prove existence and uniqueness of the charged solutions and to analyze their properties.
@@
In particular, we will show that under certain conditions, the extremal limits of the charged solutions may develop regular near–horizon regions, where the geometry reduces to a product of a sphere and AdS space, providing a multi–charge generalization of the well known examples of D3, M2, and M5 branes. Restricting to the near–horizon regime of charged geometries, one encounters not only black branes, but also regular wormhole–like solutions interpolating between two copies of the AdS space at two spacial infinities. Some solutions of this type were found in [19], and in this article we extend this construction to geometries charged under an arbitrary number of gauge fields. Although studies of wormholes have a long history [20, 21], and they intensified in the last decade [22], most results have been obtained in two and four dimensions [23, 24, 25]. Our analysis complements this work by studying higher–dimensional wormholes with multiple charges in the presence of a cosmological constant or a nontrivial potential for scalars [^6] See also [26] for other approaches to wormholes in $d>4$.. Once the cosmological parameter is switched off, the wormholes interpolating between two copies of the AdS space are no longer possible, but in this case, our geometries for multi–charged black branes become more explicit, and they extend the classic results obtained by Horowitz and Strominger [1] to an arbitrary number of scalars and higher form gauge fields.
@@ 2 Multi–charged black holes with cosmological constant
Dimensional reduction of string theory produces gravitational systems containing multiple gauge fields and scalars, and black holes charged under these fields have been extensively studied in the past [9, 7]. If the reduction is performed on a torus, then the action has only kinetic terms, and the multi–charged black holes can be constructed using algebraic techniques introduced in [7]. Reduction on spheres leads to gauged supergravities [27, 28, 29, 30], and black solutions for such systems have been found only on the case–by–case basis [31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42]. Some of these solutions are summarized in the Appendix A, where we focused on the static case studied in this article [^7] In that appendix we also give references to articles containing some rotating solutions.. In contrast to the toric reduction that generically leads to asymptotically–flat geometries, the black holes coming from gauged supergravities have AdS asymptotics, making them important for exploration of the AdS/CFT correspondence [37, 38]. In this section we will construct multi–charged black hole geometries which go beyond gauged supergravities. Specifically, we consider a system of an arbitrary number $k$ of gauge fields coupled to a correlated number of scalars and construct asymptotically–AdS black holes with $k$ charges. We will show that existence of simple closed–form solutions implies the unique potential for the scalar fields [^8] A similar logic in the case of two vector fields and one scalar was used in [43]., which generalizes results from all gauged supergravities summarized in the Appendix A. We begin with a brief summary of the setup, then our discussion will be split into two subsections: in section 2.1 we will focus on special couplings between scalars and gauge fields inspired by string theory, and in section 2.2 the black hole solutions will be extended to general couplings.
@@
$S=\int d^{d}x\sqrt{-g}\left[R-\frac{1}{2}\sum_{j=1}^{k^{\prime}}(\partial{\phi}_{j})^{2}-\sum_{I=1}^{k}\frac{1}{2(X_{I})^{2}}F^{I}\wedge\star F^{I}-{\tilde{V}}(\phi)\right]\,.$ (2.1)
@@
$S=\int dx\sqrt{-g}\left[R-w_{1}\left[\sum\partial\psi_{I}\right]^{2}-w_{2}\sum(\partial{\psi}_{I})^{2}-\sum\frac{e^{2\psi_{I}}}{2}F^{I}\wedge\star F^{I}-{V}(\psi)\right].$ (2.4)
@@ 2.1 Special case: solutions inspired by string theory
To view the geometry (2.1) as special case of static solutions for the action (2.4), we recall the standard rules for performing the dimensional reduction of the string action from $(d+1)$ to $d$ dimensions [44, 7]. For the ansatz
@@
Note that for $d=4,5$ the solution (2.6) and its rotating version were obtained in [7] by a slightly different method of compactification to three dimensions and dualizing vector fields into scalars. The combination of boosts and dualities that led to (2.6) is an equivalent construction, but since it does not rely on dualization of vectors, it works in all dimensions.
@@ 2.2 Generalization: arbitrary scalar couplings
$S=\int dx\sqrt{-g}\left[R-{\tilde{w}}_{1}\left[\sum\partial\psi_{I}\right]^{2}-{\tilde{w}}_{2}\sum(\partial{\psi}_{I})^{2}-\sum\frac{e^{\gamma\psi_{I}}}{2}F^{I}\wedge\star F^{I}-{V}(\psi)\right].$ (2.20)
@@
In gauged supergravities, some rotating extensions of (2.2) have been found as well [32, 34, 36, 39, 41, 42], but unfortunately they often require actions that go beyond (2.20), and the additional terms have been constructed only in the case–by–case basis. For example, the action of ${\cal N}=2$ gauged supergravity in five dimensions is given by (A.7),
@@
which reduces to a special case of (2.20) only in the absence of the Chern–Simons term. This term doesn’t contribute to equations of motion for the static solutions which we are considering in this article, but it is crucial for finding the rotating solutions constructed in [41]. Similar Chern–Simons couplings appear in other dimensions as well, and it would be interesting to do a reverse engineering of these terms using the logic we have implemented in this article: just as we have derived the most general potential (2.25) starting from a particular ansatz (2.2) for the static solution, it might be possible to find the higher dimensional extensions of (2.38) starting from a stationary ansatz, without appealing to gauged supergravity. An interesting example of this approach was implemented in [45], where the most general rotating solution for one gauge field was constructed in all dimensions. In this case, the Chern–Simons term never contributes, as can be seen already in the example (2.38). We hope to analyze the problem of rotating solutions with several fields in the future.
@@ 3 Multi–charged branes with scalar couplings
In the last section we constructed geometries produced by black holes charged under multiple gauge fields and scalars, and it is natural to extend this construction to black branes charged under multiple higher form fluxes. Many solutions of this type have been constructed starting from the seminal work of Horowitz and Strominger [1], and such geometries have played important role in exploration of the gauge/gravity duality. In this section we will construct solutions describing branes with multiple charges and analyze near–horizon limits of extremal geometries and conditions under which such limits reduce to products of spheres and AdS spaces.
Specifically, in section 3.1 we will use a harmonic ansatz inspired by black holes to construct black branes coupled to an arbitrary number of $p$–form fluxes. The resulting geometries have regular horizons, but in the extremal limit they may develop naked singularities. In section 3.2 we analyze such extremal solutions and determine the constraints on the number of fields which ensure non–singular behavior. In such a situation, the near–horizon limit produces a product of a sphere and an AdS space on the Poincare patch. In the same subsection we also analyze closely related wormhole–type geometries which interpolate between two copies of AdS space at two disconnected infinities. Finally, in section 3.4 we relax the harmonic ansatz and construct the most general local solutions supported by $p$–forms and scalars consistent with symmetries of black branes. The explicit solutions are more complicated than the ones obtained within the harmonic ansatz, they have additional free parameters, and generically they lead curvature singularities at the locations of the horizons.
To extend the results of the previous section to branes, we consider a generalization of the the action (2.20)
$S=\int dx\sqrt{-g}\left[R-{\tilde{w}}_{1}\left[\sum_{I}^{k}\partial\psi_{I}\right]^{2}-{\tilde{w}}_{2}\sum_{I}^{k}(\partial{\psi}_{I})^{2}-\sum_{I}^{k}\frac{e^{\gamma\psi_{I}}}{2}F_{(p+2)}^{I}\wedge\star F_{(p+2)}^{I}\right],$ (3.1)
where $F_{(p+2)}$ are $(p+2)$–form field strenghts. Note that, in contrast to (2.20), the potential $V(\psi)$ has been removed. As we have seen in the previous section, this potential is proportional to the cosmological constant, and in the presence of this parameter, there are no known analytical solutions even for neutral branes [14, 15, 18]. Therefore, we don’t expect to have simple solutions even in the charged case [^9] Note, however, that one can construct charged wormholes [19]., but some qualitative features of the resulting geometries will be analyzed in section 4.

### v2

@@ 1 Introduction
Over the last three decades, studies of black holes and black branes have led to remarkable progress in understanding string theory [1, 2, 3, 4, 5, 6], quantum gravity [7, 8, 9, 10] and strongly coupled systems [11, 12, 13, 14, 15]. In particular, an important role has been played by charged objects that generalize the classic cases of the Reissner–Nordström and Kerr–Newman black holes in several directions: they contain extended sources [1], multiple gauge fields [16, 17, 18, 19, 20, 21], nontrivial scalars [22, 23, 24], and the cosmological constant [25, 26]. While many systems without the cosmological constant arise from the ten–dimensional supergravities and their reductions, one can also extend these constructions to an arbitrary number of dimensions by starting from the Myers–Perry black holes [27] and dressing them with charges using techniques developed in [16, 17, 18, 19, 20, 21]. In the presence of the cosmological constant, the neutral solutions have also been constructed in arbitrary dimensions [25, 26], but addition of charge presents significant challenges, and has only been done on the case–by–case basis. Such charged extensions were mostly constructed in the context of gauged supergravities, which are defined only in some dimensions, and whose field contents vary significantly as one looks at different dimensions. In this article we will use known solutions of several gauged supergravities as inspirations for constructing asymptotically-AdS black holes with an arbitrary number of charges in an arbitrary number of dimensions. We will also construct closely related geometries produced by $p$–branes, extending the well-known results obtained in [1].
@@
The great success in constructing charged black holes and black branes in string theory is partly attributed to the discovery of the U–duality group that allows one to start with a neutral seed solution of nonlinear differential equations and dress it with charges using a pure algebraic procedure [^3] The discovery of black branes in supergravity by Horowitz and Strominger [1] pre-dates this algebraic method. Furthermore, the algebraic procedure applies only for specific couplings between gauge fields and scalars, and some solutions found in [1] and related work [24, 22, 23] come from more general Lagrangians.. At the level of classical gravity for bosonic fields, such U–duality group can be formally defined in any number of dimensions, allowing one to start from a product of the neutral Myers–Perry geometry [27] and a torus and add several charges to the solutions. In the process of doing so, one generates gauge fields as well as several scalars. Since the reduction is performed on a torus, the gauge fields are abelian, and scalars do not have a potential. A similar reduction on spheres is more complicated, and often it leads to gauged supergravities with a nontrivial potential for the scalars. Unfortunately, such a reduction does not produce a U–duality group large enough to algebraically generate new charged solutions from the neutral ones, so the relevant configurations have been found on the case–by–case basis by solving the equations of motion. Furthermore, the actions for gauged supergravities and even their field contents strongly depend on dimensions. On the other hand, reduction on spheres also has a significant advantage: due to a nontrivial potential, the reduced theory admits solutions with anti–de–Sitter asymptotics, while the geometries obtained from the toric case are generally asymptotically–flat [^4] One can also construct solutions with other asymptotics by taking various limits, but if all charges are turned off, then the geometries obtained via toric reduction always approach flat space at infinity.. Therefore, solutions constructed in gauged supergravities have played a prominent role in the exploration of the AdS/CFT correspondence [12, 13, 14, 15], and it is very interesting to find new examples of such geometries. In this article we will use solutions of gauged supergravities as inspirations for a general ansatz describing asymptotically AdS black holes with an arbitrary number of charges and find the unique action, which depends only on the numbers of fields and dimensions, that gives rise to such solutions. All known gauged supergravities are reproduced as special cases.
@@
While many black hole geometries with AdS asymptotics are known, finding black branes with such asymptotics presents a significant challenge, even in the neutral case. This problem can be attributed to a relatively small amount of symmetry in the resulting configuration: while black holes are naturally embedded in the AdS space in the standard coordinates, black branes violate many symmetries of this space. Therefore, the ansatz interpolating between black branes and AdS asymptotics contains many degrees of freedom, and all attempts to obtain the solutions of equations of motion in closed analytical form have been unsuccessful [^5] This discussion focuses on static flat branes. The geometries produced by supersymmetric branes with curved worldvolumes have been constructed [28, 29, 30, 31], but they contain rotation, as expected from the general properties of such branes [32, 33].. The study of extended black objects in the presence of the cosmological constant began with exploration of black strings with AdS5 asymptotics [34], and it was later extended to black strings in arbitrary dimensions in [35, 36, 37, 38, 39, 40]. The authors of these articles extracted many important properties of black strings with cosmological constant, but unfortunately, they had to rely on numerical simulations. More recent article [41] extended these results to arbitrary $p$–branes and complemented the numerical analysis of [34, 35] by various analytical expansions, and in some special cases it reduced the problem to solutions of a relatively simple Abel equation. All these analyses were performed for neutral strings/branes, and in this article we extend them to the charged case. We will use the analytical expansions introduced in [41] to prove existence and uniqueness of the charged solutions and to analyze their properties.
@@
In particular, we will show that under certain conditions, the extremal limits of the charged solutions may develop regular near–horizon regions, where the geometry reduces to a product of a sphere and AdS space, providing a multi–charge generalization of the well known examples of D3, M2, and M5 branes. Restricting to the near–horizon regime of charged geometries, one encounters not only black branes, but also regular wormhole–like solutions interpolating between two copies of the AdS space at two spacial infinities. Some solutions of this type were found in [42], and in this article we extend this construction to geometries charged under an arbitrary number of gauge fields. Although studies of wormholes have a long history [43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56], and they intensified in the last decade [57, 58, 59, 60, 61], most results have been obtained in two and four dimensions [62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72]. Our analysis complements this work by studying higher–dimensional wormholes with multiple charges in the presence of a cosmological constant or a nontrivial potential for scalars [^6] See also [73, 74, 75, 76, 77] for other approaches to wormholes in $d>4$.. Once the cosmological parameter is switched off, the wormholes interpolating between two copies of the AdS space are no longer possible, but in this case, our geometries for multi–charged black branes become more explicit, and they extend the classic results obtained by Horowitz and Strominger [1] to an arbitrary number of scalars and higher form gauge fields.
@@ 2 Multi–charged black holes with cosmological constant
Dimensional reduction of string theory produces gravitational systems containing multiple gauge fields and scalars, and black holes charged under these fields have been extensively studied in the past [24, 16, 17, 18, 19, 20, 21]. If the reduction is performed on a torus, then the action has only kinetic terms for the matter fields, and the multi–charged black holes can be constructed using algebraic techniques introduced in [16, 17, 18, 19, 20, 21]. Reduction on spheres leads to gauged supergravities [78, 79, 80, 81, 82, 83, 84], and black hole solutions for such systems have been found only on the case–by–case basis [85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96, 97, 98, 99, 100, 101, 102]. Some of these solutions are summarized in the Appendix A, where we focused on the static case studied in this article [^7] In that appendix we also give references to articles containing some rotating solutions.. In contrast to the toric reduction that generically leads to asymptotically–flat geometries, the black holes coming from gauged supergravities have AdS asymptotics, making them important for exploration of the AdS/CFT correspondence [96, 97, 98]. In this section we will construct multi–charged black hole geometries which go beyond gauged supergravities. Specifically, we consider a system of an arbitrary number $k$ of gauge fields coupled to a correlated number of scalars and construct asymptotically–AdS black holes with $k$ charges. We will show that existence of simple closed–form solutions implies the unique potential for the scalar fields [^8] A similar logic in the case of two vector fields and one scalar was used in [103]., which generalizes results from all gauged supergravities summarized in the Appendix A. We begin with a brief summary of the setup, then our discussion will be split into two subsections: in section 2.1 we will focus on special couplings between scalars and gauge fields inspired by string theory, and in section 2.2 the black hole solutions will be extended to general couplings.
@@
$S=\int d^{d}x\sqrt{-g}\left[R-\frac{1}{2}\sum_{j=1}^{k^{\prime}}(\partial{\phi}_{j})^{2}-\sum_{I=1}^{k}\frac{1}{4(X_{I})^{2}}(F_{\mu\nu}^{I})^{2}-{\tilde{V}}(\phi)\right]\,.$ (2.1)
@@
$S=\int d^{d}x\sqrt{-g}\left[R-w_{1}\left[\sum\partial\psi_{I}\right]^{2}-w_{2}\sum(\partial{\psi}_{I})^{2}-\sum\frac{e^{2\psi_{I}}}{4}(F_{\mu\nu}^{I})^{2}-{V}(\psi)\right].$ (2.4)
@@ 2.1 Special case: solutions inspired by string theory
To view the geometry (2.1) as special case of static solutions for the action (2.4), we recall the standard rules for performing the dimensional reduction of the string action from $(d+1)$ to $d$ dimensions [104, 16, 17, 18, 19, 20, 21]. For the ansatz
@@
Note that for $d=4,5$ the solution (2.6) and its rotating version were obtained in [16, 17, 18, 19, 20, 21] by a slightly different method of compactification to three dimensions and dualizing vector fields into scalars. The combination of boosts and dualities that led to (2.6) is an equivalent construction, but since it does not rely on dualization of vectors, it works in all dimensions.
@@ 2.2 Generalization: arbitrary scalar couplings
$S=\int d^{d}x\sqrt{-g}\left[R-{\tilde{w}}_{1}\left[\sum\partial\psi_{I}\right]^{2}-{\tilde{w}}_{2}\sum(\partial{\psi}_{I})^{2}-\sum\frac{e^{\gamma\psi_{I}}}{4}(F_{\mu\nu}^{I})^{2}-{V}(\psi)\right].$ (2.20)
@@
In gauged supergravities, some rotating extensions of (2.2) have been found as well [87, 89, 90, 91, 92, 93, 95, 99, 101, 102], but unfortunately they often require actions that go beyond (2.20), and the additional terms have been constructed only in the case–by–case basis. For example, the action of ${\cal N}=2$ gauged supergravity in five dimensions is given by (A.7),
@@
which reduces to a special case of (2.20) only in the absence of the Chern–Simons term. This term does not contribute to equations of motion for the static solutions which we are considering in this article, but it is crucial for finding the rotating solutions constructed in [101]. Similar Chern–Simons couplings appear in other dimensions as well, and it would be interesting to do a reverse engineering of these terms using the logic we have implemented in this article: just as we have derived the most general potential (2.25) starting from a particular ansatz (2.2) for the static solution, it might be possible to find the higher dimensional extensions of (2.38) starting from a stationary ansatz, without appealing to gauged supergravity. An interesting example of this approach was implemented in [105], where the most general rotating solution for one gauge field was constructed in all dimensions. In this case, the Chern–Simons term never contributes, as can be seen already in the example (2.38). We hope to analyze the problem of rotating solutions with several fields in the future.
@@ 3 Multi–charged branes with scalar couplings
In the last section we constructed geometries produced by black holes charged under multiple gauge fields and scalars, and it is natural to extend this construction to black branes charged under multiple higher form fluxes. Many solutions of this type have been constructed starting from the seminal work of Horowitz and Strominger [1], and such geometries have played an important role in exploration of the gauge/gravity duality. In this section we will construct solutions describing branes with multiple charges and analyze near–horizon limits of extremal geometries and conditions under which such limits reduce to products of spheres and AdS spaces.
Specifically, in section 3.1 we will use a harmonic ansatz inspired by black holes to construct black branes coupled to an arbitrary number of $p$–form fluxes. The resulting geometries have regular horizons, but in the extremal limit they may develop naked singularities. In section 3.2 we analyze such extremal solutions and determine the constraints on the number of fields which ensure non–singular behavior. In such a situation, the near–horizon limit produces a product of a sphere and an AdS space on the Poincare patch. In the same subsection we also analyze closely related wormhole–type geometries which interpolate between two copies of AdS space at two disconnected infinities. Finally, in section 3.4 we relax the harmonic ansatz and construct the most general local solutions supported by $p$–forms and scalars consistent with symmetries of black branes. The explicit solutions are more complicated than the ones obtained within the harmonic ansatz, they have additional free parameters, and generically they lead to curvature singularities at the locations of the horizons.
To extend the results of the previous section to branes, we consider a generalization of the action (2.20)
$S=\int d^{d}x\sqrt{-g}\left[R-{\tilde{w}}_{1}\left[\sum_{I}^{k}\partial\psi_{I}\right]^{2}-{\tilde{w}}_{2}\sum_{I}^{k}(\partial{\psi}_{I})^{2}-\sum_{I}^{k}\frac{e^{\gamma\psi_{I}}}{2(p+2)!}(F_{(p+2)}^{I})^{2}\right],$ (3.1)
where $F_{(p+2)}$ are $(p+2)$–form field strengths. Note that, in contrast to (2.20), the potential $V(\psi)$ has been removed. As we have seen in the previous section, this potential is proportional to the cosmological constant, and in the presence of this parameter, there are no known analytical solutions even for neutral branes [34, 35, 41]. Therefore, we do not expect to have simple solutions even in the charged case [^9] Note, however, that one can construct charged wormholes [42]., but some qualitative features of the resulting geometries will be analyzed in section 4.